I need to analyze the mass spec. For p-nitroaniline. All major fragments lost.

media%2Fba9%2Fba9043db-23f6-4ba5-ac27-4c
Answer #1

p-nitroaniline, M.W. = 138.12 g/mol

M+ = 138, indicate the molecular ion corresponds to p-nitroaniline

M/z = 108, indicate a possible Nitroso+ benzene fragment (Ph-NO+) in which Nitrobenzene lost O and become nitrosobenzne)

M/z = 92, Indicate aniline+ fragment, loss of nitro group. another possibility is the Ph-O+ ions formation from nitrobenzene+.

Intensity of M/z (108) similar to M/z (92)

M/z = 65, C5H5+ ions (cyclopenta-2,4-dien-1-ylium) degradation of benzene ring by HCN removal from aniline+. Commonly observed in aniline, anisole etc.

M/z = 39, C3H3+ ions (cyclopropenyl), loss of two CH (acetylene removal) from C5H5+ ion